Consider the solution to the IVP y - xy = x; y(0) = 2 Find y (0) Consider the solution to the IVP y - xy = t; y(0) = 2 Fin
Answer #1

First of all recall that the differential eq? of the form dy + Ple).y = Q(x)—AH(Where place s9029 da are function of se) areNow Integrating factor (Tºf) = SP(x) dx e S-xdx 11 e - (3) Te IF e multiplying Iof both side of (1), we get - 2² *(* +(234) =2 - => y e ES ne e da put -22 all t =)-2x = = feat =- Set do dt वर -Xdx=dt => x dx=-dt sret t to -e? atc -e , where c is conUsing value of e ein (f), we get g= 3 et 3 e²-1 2 ye yox) = 3 eta 2 Differentiating Wont w 를 so y(x)= 3x ye7= 3 et 22 -0